OpenStudy (alivejeremy):

5.78 * 10^-11 Put in scientific notation

OpenStudy (alivejeremy):

Tricky

OpenStudy (alivejeremy):

@YanaSidlinskiy @Ms-Brains @mathmate @inkyvoyd

OpenStudy (alivejeremy):

Is it 0.00000000578

OpenStudy (rcp031):

scientific notation? means is a special way of writing numbers: How to Do it To figure out the power of 10, think "how many places do I move the decimal point?" When the number is 10 or greater, the decimal point has to move to the left, and the power of 10 is positive. When the number is smaller than 1, the decimal point has to move to the right, so the power of 10 is negative. Example: 0.0055 is written 5.5 × 10-3 Because 0.0055 = 5.5 × 0.001 = 5.5 × 10-3 Example: 3.2 is written 3.2 × 100 We didn't have to move the decimal point at all, so the power is 100 But it is now in Scientific Notation

OpenStudy (alivejeremy):

I know what it means kiddo. ;)

OpenStudy (callie_love_me):

So what is this question?

OpenStudy (rcp031):

5.78 * 10^-11 in terms scientific notation is 578*10^(-13).

OpenStudy (alivejeremy):

no

OpenStudy (alivejeremy):

This is the question 5.78 * 10^-11

OpenStudy (alivejeremy):

Put it in scientific notation

OpenStudy (alivejeremy):

And i'm asking is this right 0.00000000578?

OpenStudy (johnweldon1993):

Hint* Take your decimal point and move it over to the right 11 times, do you get 5.78?

OpenStudy (mathmate):

@alivejeremy If you are asking for help, people take the time to explain to you. If you already know it, there are more courteous ways of saying that than `I know what it means kiddo. ;)` In fact you'll be surprised that your response shows clearly that you don't know what a scientific notation is. I suggest you reread what @rcp031 posted, and apologize to him/her about your attitude. If you still cannot figure it out, post again, but only after reading every word @rcp031 posted.

OpenStudy (alivejeremy):

I do know what scientific notation mean.

OpenStudy (alivejeremy):

I am just confused on this question.